• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. Lhuillier

    Log entry time 00:38:00 on July20,2005

    Entry number 148671

    keyword=compton tune

    Took ~1h30 of beam time to optimize background in the compton detectors. Started with S/B=2, ended up with S/B=3.
    An efficient cure last year was the adjustment of path length using the "dog-leg" magnets. This year the settings of the dog-legs are pretty constrained by the requirement of energy dispersion in hall C so we can't easily tune them. The small changes we tried tend to show that the dog-legs setting were already optimized anyway.
    Increasing the field in the quadruple IC20 (just upstream the chicane) by 10% (set point 27333-->30333) reduced the background in the photon detector by more than a factor 2. The electron rate went up by 40%. There may be room there for optimization. To be discussed tomorrow with Jay Benesh.
    The beam axis seems to be 2mm to far from the electron detector. It can be easily corrected if confirmed by further analysis.